# Explore then Determine: A GNN-LLM Synergy Framework for Reasoning over Knowledge Graph

### v1

@@ Appendix A Discussion
Limitations. In our method, the exploration depth $L$ is preset based on the characteristics of the dataset. Therefore, in real-world applications, there is a possibility that the correct answer might not be found within step $L$. It is worth exploring how to use LLM to determine the depth of exploration in the future.
Broader Impacts. The method could facilitate real-world applications, such as mentioned in Section 1. The Explore-then-Determine framework not only improves the accuracy and efficiency of KGQA tasks, but also has the potential to be adapted for a wide range of reasoning tasks involving LLMs.
Safeguards Statement. In this paper, we primarily focus on the KGQA task. We recognize that future research on border applications of reasoning with LLMs may carry the risk of misuse. Therefore, we recommend thoroughly considering all aspects of safety before applying these techniques in real-world scenarios.
@@ B.1 Question and Relation Encoding
Considering the remarkable modeling capacity of the LLM, we first employ the average embedding in the last layer of the LLM (e.g., Llama2-13B) to produce text encoding as the representations of question $q$ and relation $r$ in the KG:
@@ B.2 Exploration Algorithm
Algorithm 1 Semantic-aware graph exploration.
@@
4:     calculate attention weights $\alpha_{q|sr}^{\ell}$ (by Eq.(1));
@@
7:        $\bm{h}_{o}^{\ell}:=\delta(\sum\nolimits_{(e_{s},r,e_{o})\in{\widetilde{\mathcal{N}}_{e_{o}}^{\ell}}}\alpha_{q|sr}^{\ell}\bm{W}^{\ell}(\bm{h}_{s}^{\ell-1}\odot\bm{h}_{r}))$ (by Eq.(2)).
@@ B.5 Baselines
In Table 9, we summarize the differences between our method EtD and several representative baselines in KGQA. These methods generally involve knowledge retrieval and answer generation processes, and our exploration stage can be seen as a form of knowledge retrieval. As can be seen, our method uniquely synergizes LLM with GNN, harnessing both the precision of compositional learning and the prowess of language understanding.
@@ Appendix C Supplementary Experiments
C.1 Effect of Pre-training Strategy
We compared the performance of exploration module with pre-training and fine-tuning with those trained solely on individual datasets. We plot the learning curves for both training strategies. As can be seen in Figure 5, pre-training significantly accelerates the convergence speed of the model during fine-tuning on the downstream datasets, and acquires better performance. This advantage is more pronounced on more complex datasets (i.e., MetaQA-3hop).
@@ C.1 Effect of Pre-training Strategy
44 97.4 98.6 96.0
EtD-Llama2-13B 100 100 100
C.2 Influence of Number of Reference Answer
@@
Table 11: Comparison of different numbers of reference answers $N$ of EtD-Llama2-13B.
@@ Appendix D Case Study
Table 12 and 13 present examples of the input-output of our method EtD. Compared with Figure 4, we present the complete set of prompts, including the task description. It can be observed that by combining explicit knowledge from the KG with the implicit knowledge of LLM itself, LLM ultimately selects the correct answer and provides its reasoning process, enhancing the credibility of the output.
@@
Output by EtD-Llama2-13B Based on the reference answers and my own knowledge, I can answer the question. The correct answer is A. 1989.
@@
Output by EtD-Llama2-13B Based on the information provided, the correct answer is B. Miller Park.
@@ Appendix E Variants of Prompt
Table 14 shows the variants of designed prompt introduced in Section 4.3.3, Compared with these formats, our designed knowledge-enhanced multiple-choice prompt can achieve better performance.
@@
EtD-w.o.-mcp Reference answers include: [Seattle Pilots, Miller Park, Milwaukee].
@@
EtD-w.o.-cand Relevant facts include:
@@
EtD-w.o.-prob Reference answers:
@@
EtD-w.o.-path Reference answers:

### v2

@@ Appendix A Discussion
In our problem setting, the answers are restricted to entities within the KG, which might be too strict in real-world applications. However, as long as the answer can be linked to entities within the KG, we can still obtain supervision signals to train the GNN model and complete the entire process. Specifically, we can also leverage LLMs to generate natural questions based on the multi-hop relations between entities, and incorporating supervision signals from intermediate entities to further enhance model training. This presents a promising direction for future research.
@@ B.1 Question and Relation Encoding
Considering the remarkable modeling capacity of the LLM, we first employ the average embedding in the first and last layer of the LLM (i.e., Llama2-13B) to produce text encoding as the representations of question $q$ and relation $r$ in the KG:
@@ B.2 Exploration Algorithm
Algorithm 1 Semantic-aware knowledge exploration.
@@
4:     calculate attention weights $\alpha_{q|sr}^{\ell}$ (by (1));
@@
7:        $\bm{h}_{o}^{\ell}:=\delta(\sum\nolimits_{(e_{s},r,e_{o})\in{\widetilde{\mathcal{N}}_{e_{o}}^{\ell}}}\alpha_{q|sr}^{\ell}\bm{W}^{\ell}(\bm{h}_{s}^{\ell-1}\odot\bm{h}_{r}))$ (by (2)).
@@ B.5 Baselines
  • GNN-RAG [39] utilizes an off-the-shelf GNN model to retrieve the shortest path to high-scoring candidate entities for LLMs to reason for answers.
In Table 9, we summarize the differences between our method DualR and several representative baselines in KGQA. These methods generally involve knowledge retrieval and answer generation processes, and our knowledge exploration process can be seen as a form of knowledge retrieval. As can be seen, our method uniquely synergizes LLM with GNN, harnessing both the precision of graph learning and the prowess of language understanding.
Another category of methods, such as QA-GNN[53] and GreaseLM[54], integrates the GNN with the language model at the architectural level, making them incompatible with current LLMs. Additionally, their approach is specifically designed for multiple-choice question answering, where the model input requires candidate answers, and they lack the capability to search for candidates over the entire knowledge graph. Therefore, they are unable to complete the task in this paper.
@@ Appendix C Supplementary Experiments
C.1 Influence of Pruning
In this section, we discuss the influence of the pruning for candidate set. As shown in Figure 4, the size of the unfiltered candidate set is typically vast due to the large-scale knowledge graph. In contrast, after pruning, the candidate set is significantly reduced, which greatly decreases the computational cost during the GNN propagation process. Therefore, we can filter out a large amount of irrelevant information, significantly shortening inference time and achieving more efficient reasoning.
(a) Number of candidates.
(b) Inference time.
Figure 4: Influence of pruning on three datasets.
@@ C.1 Influence of Pruning
C.2 Effect of Pre-training Strategy
Table 10: Influence of pretraining of exploration module.
Methods WebQSP CWQ MetaQA-1 MetaQA-2 MetaQA-3
w.o.-pretrain 72.9 53.4 97.8 99.0 99.4
pretrain-finetune 76.8 55.6 97.8 99.1 99.5
In Section 3.4, we have mentioned that the GNN module can be benefited from pre-training strategies. We compare the performance of individual training strategy and pretrain-finetune strategy in Table 10. It is evident that the strategy of pre-training significantly outperforms training separately on distinct datasets. This underscores the effectiveness of pre-training in enabling our exploration module to better grasp the common compositional relationships between questions and the concepts in the graphs, thereby enhancing the question-comprehension abilities empowered by LLM.
We plot the learning curves for both training strategies. As can be seen in Figure 5, pre-training significantly accelerates the convergence speed of the model during fine-tuning on the downstream datasets, and acquires better performance. This advantage is more pronounced on more complex datasets (i.e., MetaQA-3hop).
@@ C.2 Effect of Pre-training Strategy
AlAgha [57] 97.4 98.6 96.0
DualR-Llama2-13B 100 100 100
Table 12: Performance comparison of different encoders of DualR-w.o.-AD in Hits@1(%).
Encoder WebQSP CWQ
RoBERTa-large 75.2 53.9
Llama2-13B 76.8 55.6
C.3 Influence of Question and Relation Encoder
In our GNN module, it is feasible to use a small language model (e.g., RoBERTa[58]) to encode the text of questions and relations, since the GNN module is capable of learning and adapting it. However, our experiments have shown that using Llama2 for encoding produces superior results, as evidenced by the performance presented in the table 12. This improvement is attributed to Llama2’s higher dimensionality and deeper processing in encoding, which allows it to capture semantic information and model text representations more effectively.
C.4 Influence of Number of Reference Answer
@@
Table 13: Comparison of different numbers of reference answers $N$ of DualR-Llama2-13B.
@@ Appendix D Case Study
Table 14 and 15 present examples of the input-output of our method DualR. Compared with Figure 3, we present the complete set of prompts, including the task description. It can be observed that by combining explicit knowledge from the KG with the implicit knowledge of LLM itself, LLM ultimately selects the correct answer and provides its reasoning process, enhancing the credibility of the output.
@@
Output by
Llama2-13B-CoT
The films whose writers also wrote Birdy were disseminated in the following years: * The Right Stuff (1983) * The Big Red One (1980) Both of these films were written by Larry McMurtry and Daniel Taradash, who also co-wrote the screenplay for Birdy.
Input Task
Description
Given a question, and the reference answers with their correct probabilities and associated retrieved knowledge graph triples (entity, relation, entity) as related facts, you are asked to answer the question with this information and your own knowledge. If the reference answers contain the correct answer, please output the label and content of the answer; If not, please answer the question based on your own knowledge.
@@
Output by
DualR-
Llama2-13B
Based on the reference answers and my own knowledge, I can answer the question. The correct answer is A. 1989.
@@
Question-Answer What is the most current year the team who won the 2010 World Series has won another World Series? - 2014 World Series.
Output by
Llama2-13B-CoT
The team that won the 2010 World Series, the San Francisco Giants, has not won another World Series since then. Therefore, the most current year they have won a World Series is 2010.
Input Task
Description
Given a question, and the reference answers with their correct probabilities and associated retrieved knowledge graph triples (entity, relation, entity) as related facts, you are asked to answer the question with this information and your own knowledge. If the reference answers contain the correct answer, please output the label and content of the answer; If not, please answer the question based on your own knowledge.
Input Question and Reference Answers Question: What is the most current year the team who won the 2010 World Series has won another World Series?
Reference answers:
A. 2010 World Series (correct probability: 0.487) {relevant facts: (San Francisco Giants, time.participant.event, 2010 World Series), (San Francisco Giants, sports.sports_team.championships, 2010 World Series)}
B. 2014 World Series (correct probability: 0.26) {relevant facts: (San Francisco Giants, sports.sports_team.championships, 2010 World Series), (San Francisco Giants, sports.sports_team.championships, 2014 World Series)}
C. 2012 World Series (correct probability: 0.253) {relevant facts: (San Francisco Giants, time.participant.event, 2010 World Series), (San Francisco Giants, sports.sports_team.championships, 2012 World Series)}
Output by
DualR-Llama2-13B
Based on the information provided, the most current year the team who won the 2010 World Series has won another World Series is: B. 2014 World Series. The relevant facts provided in the reference answer indicates that San Francisco Giants won 2010 World Series and 2014 World Series. Therefore, the answer is B. 2014 World Series.
Table 16: Case study 3 from CWQ.
@@
Output by DualR-Llama2-13B Based on the information provided, the correct answer is B. Miller Park.
@@ Appendix E Variants of Prompt
Table 17 shows the variants of designed prompt introduced in Section 4.3.3, Compared with these formats, our designed knowledge-enhanced multiple-choice prompt (in Table 16) can achieve better performance.
@@
DualR-w.o.-mcp Reference answers include: [Seattle Pilots, Miller Park, Milwaukee].
@@
DualR-w.o.-cand Relevant facts include:
@@
DualR-w.o.-prob Reference answers:
@@
DualR-w.o.-chain Reference answers: